. How much energy does it take to boil 750 g water from 20°C to steam at 110°C
1
Expert's answer
2020-05-29T10:04:09-0400

Let us calculate the amount of energy to heat water from 20°C to 100°C:

"Q_1=c_wm_w\Delta T_1 = 4184\,\mathrm{J/kg/K}\cdot 0.750\,\mathrm{kg}\cdot 80\,\mathrm{K} = 251040\,\mathrm{J}."

Then we should transform this water into steam, so

"Q_2 = L_wm_w = 2260000\,\mathrm{J/kg}\cdot 0.750\,\mathrm{kg} = 1695000\,\mathrm{J}."

Next, this steam should be heated, therefore

"Q_3 = c_sm_w\Delta T_2 = 2078\,\mathrm{J/kg/K}\cdot0.750\,\mathrm{kg}\cdot10\,\mathrm{K } = 15585\,\mathrm{J}."

Therefore, the total amount of energy will be

"Q_1+Q_2+Q_3 = 251040\,\mathrm{J} + 1695000\,\mathrm{J} + 15585\,\mathrm{J} = 1961625\,\mathrm{J} \approx 1.96\cdot10^6\,\mathrm{J}."